### v1

Abstract
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$. In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute (1) the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ and (2) the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for all $1\leq j\leq i$. We propose an $O(n)$-space data structure which computes (1) and (2) in $O(|S_{i}|\log\sigma+i)$ time for each new given string $S_{i}$, where $n$ is the total length of the strings.
Keywords:
all-pairs suffix-prefix problem Aho-Corasick automata directed acyclic word graphs (DAWGs)
1 Introduction
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics, since it is the first step of genome assembly [9]. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$.
A straightforward solution, mentioned in [10], is to use a variant of the KMP pattern matching algorithm [13] for each pair of two strings $S_{i},S_{j}$ in $O(|S_{i}|+|S_{j}|)$ time. This however leads to an inefficient $O(kn)$-time complexity, where $n=\|\mathcal{S}\|$ is the total length of the strings in $\mathcal{S}$.
Gusfield et al. [10] proposed the first efficient solution to the APSP problem that takes $O(n)$ space, which is based on the generalized suffix tree [5, 10] for the set $\mathcal{S}$ of strings. After building the generalized suffix tree in $O(n\log\sigma)$ time in the case of a general ordered alphabet of size $\sigma$ [22] or in $O(n)$ time in the case of an integer alphabets of size $\sigma=n^{O(1)}$ [8], Gusfield et al.’s algorithm [10] works in $O(n+k^{2})$ optimal time. Ohlebusch and Gog [18] proposed an alternative algorithm for solving the APSP problem with the same complexities as above, using (enhanced) suffix arrays [17, 1]. Tustumi et a. [21] gave an improved suffix-array based algorithm that is fast and memory efficient in practice. All these approaches share the common concepts of building the generalized suffix tree/array for $\mathcal{S}$.
Another data structure that can be used to solve the APSP problem is the Aho-Corasick automata (AC-automata[2]. This is intuitive since the AC-automaton is a generalization of the KMP-automaton for multiple strings. The use of the AC-automaton for solving APSP was suggested by Lim and Park [14], but they did not give any details of an algorithm nor the complexity. Recently, Loukides and Pissis [15] proposed an AC-automaton based algorithm for the APSP problem. After building the AC-automaton in $O(n\log\sigma)$ time in the case of a general ordered alphabet of size $\sigma$ [2] or in $O(n)$ time in the case of an integer alphabets of size $\sigma=n^{O(1)}$ [7], the algorithm of Loukides and Pissis [15] runs in optimal $O(n+k^{2})$ time. Their algorithm can also solve a length-threshold version of the problem in optimal time. Recently, Loukides et al. [16] considered a query-version of the APSP problem. They presented a data structure of $O(n)$-space that can report the longest suffix-prefix match between one string $S_{i}$ and all the other strings in $O(k)$ time. Their data structure is based on the AC-automaton and the micro-macro decomposition of trees [3].
In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i}=\{S_{1},\ldots,S_{i-1}\}$ of strings, to compute the following longest suffix-prefix matches between a new string $S_{i}$ and all the other strings $S_{1},\ldots,S_{i-1}$:
  • the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for $1\leq j\leq i$, and
  • the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for $1\leq j\leq i$.
We propose an $O(n)$-space data structure based on the directed acyclic word graph (DAWG) for a growing set of strings, which is able to compute the suffix-prefix matches for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time. After iterating this for all $k$ strings $S_{1},\ldots,S_{k}$ arriving to $\mathcal{S}$, the total running time becomes $O(n\log\sigma+k^{2})$, which is as fast as the state-of-the-art algorithms [10, 15] for a static set of strings in the case of general ordered alphabets of size $\sigma$. We note that the suffix-tree based algorithm of Gusfield et al. [10] performs a DFS, and thus, does not seem to extend directly to the dynamic case.
Before describing our APSP algorithm for the dynamic case (Section 5), we present a new AC-automata based APSP algorithm for a static set of strings (Section 4). While the algorithm of Loukides and Pissis [15] is based on a reduction of the APSP problem to the ULIT (union of labeled intervals on a tree) problem, our new APSP algorithm does not use this reduction. Instead, we use simple failure link traversal of the AC-automaton and marking operations on a compact prefix trie that is a compact version of the AC-trie. The simple nature of our algorithm makes it possible to work also on the DAWG structure in the dynamic case.
The hierarchical overlap graphs (HOGs) and their variants have a close relationship to the APSP problem. There are AC-automata oriented approaches for computing these graphs [6, 12, 19]. We remark that all of their approaches are for the static case, and do not seem to immediately extend to the dynamic setting.
2 Preliminaries
2.1 Strings
Let $\Sigma$ denote an ordered alphabet of size $\sigma$. An element of $\Sigma^{*}$ is called a string. The length of a string $S\in\Sigma^{*}$ is denoted by $|S|$. The empty string $\varepsilon$ is the string of length 00. Let $\Sigma^{+}=\Sigma^{*}\setminus\{\varepsilon\}$. For string $S=xyz$, $x$, $y$, and $z$ are called the prefix, substring, and suffix of $S$, respectively. Let $\mathsf{Prefix}(S)$, $\mathsf{Substr}(S)$, and $\mathsf{Suffix}(S)$ denote the sets of prefixes, substrings, and suffixes of $S$, respectively. The elements of $\mathsf{Prefix}(S)\setminus\{S\}$, $\mathsf{Substr}(S)\setminus\{S\}$, and $\mathsf{Suffix}(S)\setminus\{S\}$ are called the proper prefixes, proper substrings, and proper suffixes of $S$, respectively. For a string $S$ of length $n$, $S[i]$ denotes the $i$-th symbol of $S$ and $S[i..j]=S[i]\cdots S[j]$ denotes the substring of $S$ that begins at position $i$ and ends at position $j$ for $1\leq i\leq j\leq n$. For convenience, let $S[i..j]=\varepsilon$ for $i>j$. The reversed string of a string $S$ is denoted by $S^{R}$, that is, $S^{R}=S[|S|]\cdots S[1]$.
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, let $\mathsf{Substr}(\mathcal{S})=\bigcup_{i=1}^{k}\mathsf{Substr}(S_{i})$.
2.2 All-Pairs Suffix-Prefix Overlap (APSP) Problems
In this section, we introduce the problems that we tackle in this paper.
Let $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ be a set of non-empty strings, where $S_{i}\in\Sigma^{+}$ for $1\leq i\leq k$. For each ordered pair $\langle S_{i},S_{j}\rangle$ of two strings in $\mathcal{S}$, the longest string in the set $\mathsf{Suffix}(S_{i})\cap\mathsf{Prefix}(S_{j})$ is called the longest suffix-prefix overlap of $S_{i}$ and $S_{j}$, and is denoted by $\mathsf{lspo}(i,j)$. We encode each $\mathsf{lspo}(i,j)$ by a tuple $(i,j,|\mathsf{lspo}(i,j)|)$ using $O(1)$ space.
2.2.1 APSP Problems on Static Sets.
Problem 1 (static-APSP)
The static all-pairs suffix-prefix overlaps (static-APSP) problem is, given a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ non-empty strings, to compute
$\mathcal{L_{S}}=\{\mathsf{lspo}(i,j)\mid 1\leq i\leq k,1\leq j\leq k\}.$
Note that the output size of Problem 1 is $|\mathcal{L_{S}}|=\Theta(k^{2})$ as we compute $\mathsf{lspo}(i,j)$ for all ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$.
Problem 2 (length-bounded static-APSP)
The length-bounded static-APSP problem is, given a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ non-empty strings and an integer threshold $\ell\geq 0$, compute
$\mathcal{L_{S}^{\geq\ell}}=\{\mathsf{lspo}(i,j)\mid 1\leq i\leq k,1\leq j\leq k,|\mathsf{lspo}(i,j)|\geq\ell\}.$
2.2.2 APSP Problems on Dynamic Sets with Insertions.
Let $\mathcal{S}_{0}=\emptyset$ and for $1\leq i\leq k$, let $\mathcal{S}_{i}=\{S_{1},\ldots,S_{i}\}$ denote the subset of $\mathcal{S}$ consisting of the first $i$ strings. We consider a dynamic variant of APSP allowing for insertions of strings, named dynamic-APSP, that is defined as follows:
Problem 3 (dynamic-APSP)
The dynamic-APSP problem is, each time a new string $S_{i}$ is added to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$, to compute the two following sets
$\mathcal{F}_{i}$ $=$ $\{\mathsf{lspo}(i,j)\mid 1\leq j\leq i\},$
$\mathcal{B}_{i}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i\}.$
We remark that $\mathcal{L_{S}}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}\cup\mathcal{B}_{i}\right)$ holds for $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$. Thus, a solution to Problem 3 is also a solution to Problem 1.
Problem 4 (length-bounded dynamic-APSP)
Let $\ell>0$ be a length threshold. The dynamic-APSP problem is, each time a new string $S_{i}$ is added to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$, to compute the two following sets
$\mathcal{F}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(i,j)\mid 1\leq j\leq i,|\mathsf{lspo}(i,j)|\geq\ell\},$
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(i,j)|\geq\ell\}.$
Note that a solution to Problem 4 is also a solution to Problem 2 as well.
3 Tools
In this section we list the data structures that are used as building blocks of our static-APSP and dynamic-APSP algorithms.
3.1 Tries and Compact Tries
A trie $\mathsf{T}$ is a rooted tree $(\mathsf{V},\mathsf{E})$ such that
  • each edge in $\mathsf{E}$ is labeled by a single character from $\Sigma$ and
  • the character labels of the out-going edges of each node begin with distinct characters.
For a pair of nodes $u,v$ in a trie where $u$ is an ancestor of $v$, let $\langle u,v\rangle$ denote the path from $u$ to $v$. The path label of $\langle u,v\rangle$ is the concatenation of the edge labels from $u$ to $v$. The string label of a node $v$ in a trie, denoted $\mathsf{str}(v)$, is the path label from the root to $v$. The string depth of a node $v$ in a trie is the length of the string label of $v$, that is $|\mathsf{str}(v)|$.
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, let $\mathsf{Trie}(\mathcal{S})$ denote the trie that represents the strings in $\mathcal{S}$. We assume that each of the $k$ nodes in $\mathsf{Trie}(\mathcal{S})$ representing $S_{i}$ stores a unique integer $i$ called the id.
We build and maintain a compacted trie $\mathsf{ComTrie}(\mathcal{S})$ that is obtained from $\mathsf{Trie}(\mathcal{S})$ by keeping
  • all the leaves and the (possibly non-branching) internal nodes that store id’s $1,\ldots,k$ for the $k$ strings $S_{1},\ldots,S_{k}$, and
  • the root and all the branching nodes.
All other other nodes, each having a single child, are removed and their corresponding edges are contracted into paths. For each node $v$ in $\mathsf{Trie}(\mathcal{S})$, let $\mathsf{des}(v)$ denote the shallowest descendant of $v$ that exists in $\mathsf{ComTrie}(\mathcal{S})$. Note that the number of nodes in $\mathsf{ComTrie}(\mathcal{S})$ is $O(k)$.
Figure 1: Illustrations of $\mathsf{AC}(\mathcal{S})$ (left) and $\mathsf{ComTrie}(\mathcal{S})$ (right) for the set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings. The bold solid arcs represent trie edges and the dashed arcs represent failure links. The nodes representing the strings in $\mathcal{S}$ are depicted by double-lined circles with the string id’s.
3.2 Aho-Corasick Automata
For each non-root node $v$ of $\mathsf{Trie}(\mathcal{S})$, define its failure link by $\mathsf{flink}(v)=u$ iff $\mathsf{str}(u)$ is the longest proper suffix of $\mathsf{str}(v)$ that can be spelled out from the root of $\mathsf{Trie}(\mathcal{S})$. In other words, $\mathsf{str}(u)$ is the longest prefix of some string(s) represented by $\mathsf{Trie}(\mathcal{S})$ that is also a proper suffix of $\mathsf{str}(v)$. The failure links also form a (reversed) tree, and let us denote it by $\mathsf{FLTree}(\mathcal{S})$.
The Aho-Corasick automaton [2] for a set $\mathcal{S}$ of strings, denoted $\mathsf{AC}(\mathcal{S})$, is a finite automaton with two kinds of transitions [^1] We do not use the output function in our algorithms.: the goto function that is represented by the edges of the trie $\mathsf{Trie}(\mathcal{S})$, and the failure function that is represented by the edges of the reversed trie $\mathsf{FLTree}(\mathcal{S})$. It is clear that the total number of nodes and edges of $\mathsf{AC}(\mathcal{S})$ is $O(n)$, where $n=\|\mathcal{S}\|$. See Fig. 1 for concrete examples of the AC-automaton and its trie, compact trie, and failure links.
Theorem 3.1 ([7, 2])
For a set $\mathcal{S}$ of strings of total length $n$, $\mathsf{AC}(\mathcal{S})$ can be built
  • in $O(n)$ time and space for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma)$ time and $O(n)$ space for a general ordered alphabet of size $\sigma$.
3.3 Directed acyclic word graphs (DAWGs)
For a set $\mathcal{S}$ of strings, let $\mathsf{End\_Pos}_{\mathcal{S}}(w)$ denote the set of pairs of string ids and ending positions of all occurrences of a substring $w\in\mathsf{Substr}(\mathcal{S})$, that is,
$\mathsf{End\_Pos}_{\mathcal{S}}(w)=\{(i,j)\mid S_{i}[j-|w|+1..j]=w,1\leq i\leq k,1\leq j\leq|S_{i}|\}.$
We consider an equivalence relation $\equiv_{\mathcal{S}}$ of strings over $\Sigma$ w.r.t. $\mathcal{S}$ such that, for any two strings $w$ and $u$, $w\equiv_{\mathcal{S}}u$ iff $\mathsf{End\_Pos}_{\mathcal{S}}(w)=\mathsf{End\_Pos}_{\mathcal{S}}(u)$. For any string $x\in\Sigma^{*}$, let $[x]_{\mathcal{S}}$ denote the equivalence class for $x$ w.r.t. $\equiv_{\mathcal{S}}$.
Figure 2: Illustration of $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab}$,
$\mathrm{babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is $\mathsf{Trie}(\mathcal{S})$.
Figure 3: Illustration of the suffix links of $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}$ of strings as in Fig 2.
The DAWG of a set $\mathcal{S}$ of strings, denoted $\mathsf{DAWG}(\mathcal{S})$, is an edge-labeled DAG $(V,E)$ such that
$V$ $=$ $\{[{x}]_{\mathcal{S}}\mid x\in\mathsf{Substr}(\mathcal{S})\},$
$E$ $=$ $\{([{x}]_{\mathcal{S}},b,[{xb}]_{\mathcal{S}})\mid x,xb\in\mathsf{Substr}(\mathcal{S}),b\in\Sigma\}.$
We also define the set $L$ of suffix links of $\mathsf{DAWG}(\mathcal{S})$ by
$L=\{([{ax}]_{\mathcal{S}},a,[{x}]_{\mathcal{S}})\mid x,ax\in\mathsf{Substr}(\mathcal{S}),a\in\Sigma,[{ax}]_{\mathcal{S}}\neq[{x}]_{\mathcal{S}}\}.$
Namely, two substrings $x$ and $y$ in $\mathsf{Substr}(\mathcal{S})$ are represented by the same node of $\mathsf{DAWG}(\mathcal{S})$ iff the ending positions of $x$ and $y$ in the strings of $\mathcal{S}$ are equal.
For a node $v\in V$ of $\mathsf{DAWG}(\mathcal{S})$, let $\mathsf{long}(v)$ denote the longest string represented by $v$ (i.e., $|\mathsf{long}(v)|$ is the length of the longest path from the source to $v$). Also, let $\mathsf{slink}(v)=u$ denote the suffix link from node $u$ to node $v$.
An edge $(u,a,v)\in E$ of $\mathsf{DAWG}(\mathcal{S})$ is called an primary edge if $|\mathsf{long}(u)|+1=|\mathsf{long}(v)|$, and it is called a secondary edge otherwise ($|\mathsf{long}(u)|+1<|\mathsf{long}(v)|$). The primary edges form a spanning tree of $\mathsf{DAWG}(\mathcal{S})$ which consists of the longest paths from the source to all the nodes, and let us denote this spanning tree by $\mathsf{LPTree}(\mathcal{S})$. By the definition of the equivalence class $[{\cdot}]_{\mathcal{S}}$, for each $S_{i}\in\mathcal{S}$ it holds that $S_{i}=\mathsf{long}([{S_{i}}]_{\mathcal{S}})$. Thus we have the following:
Observation 1
For each $S_{i}\in\mathcal{S}$, the path that spells out $S_{i}$ from the source of $\mathsf{DAWG}(\mathcal{S})$ is a path of $\mathsf{LPTree}(\mathcal{S})$.
$\mathsf{DAWG}(\mathcal{S})$ is a partial DFA of size $O(n)$ that accepts $\mathsf{Substr}(\mathcal{S})$ [4, 5]. The following lemma permits us to efficiently update the DAWG each time a new string is inserted to the current set of strings:
Theorem 3.2 ([5])
Suppose that $\mathsf{DAWG}(\mathcal{S}_{i-1})$ has been built for a set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, one can update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time.
4 Algorithm for Static-APSP
In this section, we present a new, very simple algorithm for the static-APSP problem for a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $\|\mathcal{S}\|=n$.
Our main data structure is the AC-automaton $\mathsf{AC}(\mathcal{S})$. For each string $S_{i}\in\mathcal{S}$ (in an arbitrary order), set node $v$ to be the node that represents $S_{i}$, i.e. initially $\mathsf{str}(v)=S_{i}$. All the nodes of $\mathsf{ComTrie}(\mathcal{S})$ are initially unmarked. Our algorithm is shown in Algorithm 1. See also Fig. 4 for illustration.
{itembox}
[c]Algorithm 1: Compute $\mathsf{lspo}(i,j)$ for all $1\leq j\leq k$
  1. (1)
    Set $v$ to be the node that represents $S_{i}$ (i.e. $\mathsf{str}(v)=S_{i}$).
  2. (2)
    • Set $u\leftarrow\mathsf{des}(v)$.
    • Traverse the induced subtree $\mathsf{T}_{u}$ of $\mathsf{ComTrie}(\mathcal{S})$ that is rooted at $u$, and consists only of unmarked nodes (and thus we do not traverse the subtrees under the marked nodes in $T_{u}$).
    • Report $\mathsf{lspo}(i,j)=|\mathsf{str}(v)|$ for all id’s $j$ found in the induced subtree $\mathsf{T}_{u}$.
    • Mark node $u$ on $\mathsf{ComTrie}(\mathcal{S})$.
  3. (3)
    If $v$ is the root, finish the procedure for the given string $S_{i}$. Unmark all marked nodes.
  4. (4)
    If $v$ is not the root, take the failure link from $v$ on $\mathsf{Trie}(\mathcal{S})$, and set $v\leftarrow\mathsf{flink}(v)$. Continue Step (2).
Figure 4: Illustration for Algorithm 1 that solves the static-APSP problem, in which we have climbed up the failure links from the node with id $i$ to node $v$. We access $u=\mathsf{des}(v)$ in $\mathsf{ComTrie}(\mathcal{S})$, and traverse only the white subtree below $u$.
Theorem 4.1
For a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $n$, our algorithm solves the static-APSP problem with $O(n)$ working space and
  • in $O(n+k^{2})$ time for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma+k^{2})$ time for a general alphabet of size $\sigma$.
Proof
The correctness of our algorithm follows from the definitions of the $\mathsf{lspo}(\cdot,\cdot)$ function and the failure links of $\mathsf{AC}(\mathcal{S})$, and from our failure link traversal starting from the node $v$ representing $S_{i}$, in decreasing order of the node depths towards the root $r$.
For each string $S_{i}$, Step (1) clearly takes $O(1)$ time. The total number $l$ of nodes $v,\mathsf{flink}(v),\mathsf{flink}^{2}(v),\ldots,\mathsf{flink}^{l-1}(v)=r$ visited in the chain of failure links starting from node $v$ with $\mathsf{str}(v)=S_{i}$ is at most $|S_{i}|+1$. The size of $\mathsf{ComTrie}(\mathcal{S})$ is $O(k)$. As we process the nodes $v,\mathsf{flink}(v),\mathsf{flink}^{2}(v),\ldots,\mathsf{flink}^{l-1}(v)=r$ in decreasing order of their depths, any marked node is visited at most once. This is because, if a marked node $w$ is visited twice, then there must be two distinct ancestors $p,q$ of $w$ in $\mathsf{ComTrie}(\mathcal{S})$, such that $p$ is an ancestor of $q$. This however is a contradiction, since $q$ was marked before $p$, and our algorithm does not visit the subtree under $q$ when processing the induced tree $\mathsf{T}_{p}$. Therefore, each node is marked at most once, and each marked node is visited at most once. Thus, the traversals on $\mathsf{ComTrie}(\mathcal{S})$ with our marking operations takes time linear in its size $O(k)$. Overall, for each $S_{i}\in\mathcal{S}$, we can compute $\mathsf{lspo}(i,j)$ for all $1\leq i\leq k$ in optimal $O(|S_{i}|+i)$ time, which leads to an overall $O(n+k^{2})$ time for all the $k$ strings in $\mathcal{S}$.
We build $\mathsf{AC}(\mathcal{S})$ by Theorem 3.1. We can compute $\mathsf{ComTrie}(\mathcal{S})$ and $\mathsf{des}(u)=v$ for all nodes $v$ in $\mathsf{Trie}(\mathcal{S})$ in time linear in the size of $\mathsf{Trie}(\mathcal{S})$, which is $O(n)$.
For the length-bounded APSP problem, we have the following:
Corollary 1
For a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $n$, our algorithm solves the $-\ell$ static-APSP problem with $O(n)$ working space and
  • in $O(n+|\mathcal{L_{\mathcal{S}}^{\geq\ell}}|)$ time for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma+|\mathcal{L_{\mathcal{S}}^{\geq\ell}}|)$ time for a general alphabet of size $\sigma$.
Our data structure can be reused for different values of the length-threshold $\ell$.
Proof
It suffices for us to change Step (3) and Step (4) in Algorithm 1 as follows:
  1. (3)
    If $|\mathsf{str}(v)|\leq\ell$, finish the procedure for the given string $S_{i}$. Unmark all marked nodes.
  2. (4)
    If $|\mathsf{str}(v)|>\ell$, take the failure link from $v$ on $\mathsf{Trie}(\mathcal{S})$, and set $v\leftarrow\mathsf{flink}(v)$. Continue Step (2).
We can reuse our data structure for different values of $\ell$, as we only need to change the underlined conditions in the above.
5 Algorithm for Dynamic-APSP
In this section, we present our algorithm for solving the dynamic-APSP problem: Given a new string $S_{i}$ that is inserted to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, our goal is to compute the outputs $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ of Problem 3.
In what follows, we will show the following:
Theorem 5.1
There exists a data structure of $O(n_{i})$ space, which solves the dynamic-APSP problem of computing $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
Corollary 2
Let $\ell>0$ be a length threshold. There exists a data structure of $O(n_{i})$ space, which solves the length-bounded dynamic-APSP problem of computing $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
We will focus on proving Theorem 5.1. Corollary 2 can be obtained from Theorem 5.1, in a similar manner to the static case in Section 4.
5.1 Computing $\mathcal{F}_{i}$
We use the DAWG data structure for the dynamic case. Consider a set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. By Observation 1, $\mathsf{Trie}(\mathcal{S}_{i-1})$ is an induced tree of $\mathsf{LPTree}(\mathcal{S}_{i-1})$ that consists only of the paths spelling out $S_{1},\ldots,S_{i-1}$ from the source of $\mathsf{DAWG}(\mathcal{S}_{i-1})$. Also, if $\mathsf{flink}(u)=v$ for a node $u$ of $\mathsf{Trie}(\mathcal{S}_{i-1})$, then there is a chain of suffix links $u,\mathsf{slink}(u),\ldots,v$ of length $k\geq 1$ from the node $u$ to the node $v$ in $\mathsf{DAWG}(\mathcal{S}_{i-1})$ (i.e. $\mathsf{slink}^{k}(u)=v$). It is known that the suffix links of $\mathsf{DAWG}(\mathcal{S}_{i-1})$ form an edge-reversed tree that is equivalent to the suffix tree [22] of the reversed strings [5, 20]. We denote this suffix link tree by $\mathsf{SLTree}(\mathcal{S})$, and will use $\mathsf{SLTree}(\mathcal{S})$ in the dynamic case, instead of the tree $\mathsf{FLTree}(\mathcal{S})$ of failure links in the static case.
Now, we design our algorithm for the dynamic case with the DAWG structure, using the idea from Section 4 for the static case with the AC-automaton (see also Algorithm 1). Suppose that we have built and maintained the aforementioned data structures for the dynamic set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, we update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time with Theorem 3.2. We have now updated $\mathsf{SLTree}(\mathcal{S}_{i-1})$ to $\mathsf{SLTree}(\mathcal{S}_{i})$ as well.
To perform equivalent procedures as in Algorithm 1, we climb up the suffix link path from the node $v$ with $\mathsf{str}(v)=S_{i}$ toward the root on $\mathsf{SLTree}(\mathcal{S}_{i})$. Namely, we use $\mathsf{slink}$ instead of $\mathsf{flink}$. Each time we visit a node $v$ that is contained in $\mathsf{Trie}(\mathcal{S}_{i})$, we access $u=\mathsf{des}(v)$ on $\mathsf{ComTrie}(\mathcal{S}_{i})$ and perform the same procedures as in Algorithm 1. This gives us $\mathsf{lspo}(i,j)$ for all $1\leq j\leq i$ in a total of $O(|S_{i}|\log\sigma+i)$ time, since $\mathsf{ComTrie}(\mathcal{S}_{i})$ is of size $O(i)$. We can determine whether a node $v$ is in $\mathsf{Trie}(\mathcal{S}_{i})$ or not in $O(1)$ time after a simple $O(|S_{i}|\log\sigma)$-time preprocessing per new string $S_{i}$: After inserting $S_{i}$ to the DAWG, we simply trace the path that spells out $S_{i}$ from the source on the DAWG.
What remains is how to update $\mathsf{ComTrie}(\mathcal{S}_{i-1})$ to $\mathsf{ComTrie}(\mathcal{S}_{i})$, and update the $\mathsf{des}$ function. We do this by naïvely inserting the new string $S_{i}$ to $\mathsf{ComTrie}(\mathcal{S}_{i-1})$, traversing the tree with $S_{i}$ from the root in $O(|S_{i}|\log\sigma)$ time. The $\mathsf{des}$ function needs to be updated when an edge $(v,u)$ on $\mathsf{ComTrie}(\mathcal{S}_{i-1})$ is split into two edges $(v,w)$ and $(w,u)$, where $w$ is the parent of the new leaf representing $S_{i}$ on $\mathsf{ComTrie}(\mathcal{S}_{i})$. Let $v=v_{1},\ldots,v_{x}=w$ be the nodes in the uncompacted trie $\mathsf{Trie}(\mathcal{S}_{i})$ that are implicit on the edge $(v,w)$ in the compacted trie $\mathsf{ComTrie}(\mathcal{S}_{i})$. We update $\mathsf{des}(v_{1})\leftarrow w,\ldots,\mathsf{des}(v_{x})\leftarrow w$. We set $\mathsf{des}(y)\leftarrow s$ for all new nodes $y$ of $\mathsf{Trie}(\mathcal{S}_{i})$, where $s$ is the new leaf representing $S_{i}$. These updates can be done in $O(|S_{i}|)$ time.
To summarize this subsection, we have proven the following:
Lemma 1
There is a data structure of $O(n_{i})$ space, which computes $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
Remark 1
Diptarama et al. [11] showed how to simulate the failure link tree of the AC-automaton with the suffix link tree of the DAWG that is augmented with a nearest marked ancestor (NMA) data structure [23]. It should be noted that NMA data structures are not required in our algorithm.
5.2 Computing $\mathcal{B}_{i}$
We can easily reduce the problem of computing $\mathcal{B}_{i}$ to computing $\mathcal{F}_{i}$ for the set $\mathcal{S}_{i}^{R}=\{S_{1}^{R},\ldots,S_{i}^{R}\}$ of reversed strings, where the roles of suffixes and prefixes are swapped. This immediately gives us the following:
Corollary 3
There is a data structure of $O(n_{i})$ space, which computes $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
6 Conclusions and Future Work
In this paper, we presented an efficient algorithm for the APSP problem for a dynamic set of strings $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of total length $n$. Our algorithm occupies $O(n)$ space and takes $O(|S_{i}|\log\sigma+i)$ time for each string $S_{i}$ inserted, for $1\leq i\leq k$. The running time is near-optimal except for the $\log\sigma$ factor, that is indeed required if one maintains $\mathcal{S}$ in a tree or DAG based data structure for a general ordered alphabet of size $\sigma$.
An intriguing question is whether our method can be modified to handle deletions of strings from $\mathcal{S}$. Diptarama et al. [11] showed how to update the failure links of the AC-automaton in $O(\sigma|S|+u_{f})$ time when a string $S$ is deleted from $\mathcal{S}$, where $u_{f}$ is the number of failure links that need to be updated. It is interesting to see whether our DAWG-based method can be modified for the fully dynamic set of strings allowing for both insertions and deletions.
It is also interesting to see whether our dynamic algorithm can be extended to efficient construction/update of the HOG [6, 12, 19] for a dynamic set of strings.
{credits}
6.0.1 Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers and JP20H05964, JP23K24808, JP23K18466 (SI).