### v2

# All-Pairs Suffix-Prefix on Fully Dynamic Set of Strings

Abstract
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$. In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute (1) the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ and (2) the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for all $1\leq j\leq i$. We propose an $O(n)$-space data structure which computes (1) and (2) in $O(|S_{i}|\log\sigma+i)$ time for each new given string $S_{i}$, where $n$ is the total length of the strings and $\sigma$ is the alphabet size. Further, we show how to extend our methods to the fully dynamic version of the APSP problem allowing for both insertions and deletions of strings.
1 Introduction
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics, since it is the first step of genome assembly [9]. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$.
A straightforward solution, mentioned in [10], is to use a variant of the KMP pattern matching algorithm [14] for each pair of two strings $S_{i},S_{j}$ in $O(|S_{i}|+|S_{j}|)$ time. This however leads to an inefficient $O(kn)$-time complexity, where $n=\|\mathcal{S}\|$ is the total length of the strings in $\mathcal{S}$.
Gusfield et al. [10] proposed the first efficient solution to the APSP problem that takes $O(n)$ space, which is based on the generalized suffix tree [5, 10] for the set $\mathcal{S}$ of strings. After building the generalized suffix tree in $O(n\log\sigma)$ time in the case of a general ordered alphabet of size $\sigma$ [27] or in $O(n)$ time in the case of an integer alphabets of size $\sigma=n^{O(1)}$ [8], Gusfield et al.’s algorithm [10] works in $O(n+k^{2})$ optimal time. Ohlebusch and Gog [21] proposed an alternative algorithm for solving the APSP problem with the same complexities as above, using (enhanced) suffix arrays [20, 1]. Tustumi et a. [25] gave an improved suffix-array based algorithm that is fast and memory efficient in practice. All these approaches share the common concepts of building the generalized suffix tree/array for $\mathcal{S}$.
Another data structure that can be used to solve the APSP problem is the Aho-Corasick automata (AC-automata[2]. This is intuitive since the AC-automaton is a generalization of the KMP-automaton for multiple strings. The use of the AC-automaton for solving APSP was suggested by Lim and Park [17], but they did not give any details of an algorithm nor the complexity. Recently, Loukides and Pissis [18] proposed an AC-automaton based algorithm for the APSP problem. After building the AC-automaton in $O(n\log\sigma)$ time in the case of a general ordered alphabet of size $\sigma$ [2] or in $O(n)$ time in the case of an integer alphabets of size $\sigma=n^{O(1)}$ [7], the algorithm of Loukides and Pissis [18] runs in optimal $O(n+k^{2})$ time. Their algorithm can also solve a length-threshold version of the problem in optimal time. Recently, Loukides et al. [19] considered a query-version of the APSP problem. They presented a data structure of $O(n)$-space that can report the longest suffix-prefix match between one string $S_{i}$ and all the other strings in $O(k)$ time. Their data structure is based on the AC-automaton and the micro-macro decomposition of trees [3].
In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i}=\{S_{1},\ldots,S_{i-1}\}$ of strings, to compute the following longest suffix-prefix matches between a new string $S_{i}$ and all the other strings $S_{1},\ldots,S_{i-1}$:
  • the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for $1\leq j\leq i$, and
  • the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for $1\leq j\leq i$.
We propose an $O(n)$-space data structure based on the directed acyclic word graph (DAWG) for a growing set of strings, which is able to compute the suffix-prefix matches for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time. After iterating this for all $k$ strings $S_{1},\ldots,S_{k}$ arriving to $\mathcal{S}$, the total running time becomes $O(n\log\sigma+k^{2})$, which is as fast as the state-of-the-art algorithms [10, 18] for a static set of strings in the case of general ordered alphabets of size $\sigma$. We note that the suffix-tree based algorithm of Gusfield et al. [10] performs a DFS, and thus, does not seem to extend directly to the dynamic case.
Further, we consider the fully dynamic APSP problem which allows for both insertions and deletions of strings. We show how to solve fully dynamic APSP in $O(|S_{i}|\log\sigma+k)$ amortized time when string $S_{i}$ is inserted or deleted, where $k$ is the number of strings in the current set $\mathcal{S}$ of strings. In this fully dynamic version, we show how the DAWG-based algorithm can be simulated with the suffix tree in the case of insertions, and how our suffix-tree based method can handle the case of deletions. We emphasize that our approach is again quite different from Gusfield et al.’s method [10].
Before describing our APSP algorithms for the (fully) dynamic case (Section 5 and Section 6), we present a new AC-automata based APSP algorithm for a static set of strings (Section 4). While the algorithm of Loukides and Pissis [18] is based on a reduction of the APSP problem to the ULIT (union of labeled intervals on a tree) problem, our new APSP algorithm does not use this reduction. Instead, we use simple failure link traversals of the AC-automaton, and marking operations on a compact prefix trie that is a compact version of the AC-trie. The simple nature of our algorithm makes it possible to work also on the DAWG and on the suffix tree in the (fully) dynamic case.
The hierarchical overlap graphs (HOGs) and their variants have a close relationship to the APSP problem. There are AC-automata oriented approaches for computing these graphs [6, 12, 22]. We remark that all of their approaches are for the static case, and do not seem to immediately extend to the dynamic setting.
A preliminary version of this paper appeared in [13]. The new material in this full version is the algorithm for the fully dynamic APSP problem (Section 6), which was posed as an open question in the preliminary version [13].
2 Preliminaries
2.1 Strings
Let $\Sigma$ denote an ordered alphabet of size $\sigma$. An element of $\Sigma^{*}$ is called a string. The length of a string $S\in\Sigma^{*}$ is denoted by $|S|$. The empty string $\varepsilon$ is the string of length 00. Let $\Sigma^{+}=\Sigma^{*}\setminus\{\varepsilon\}$. For string $S=xyz$, $x$, $y$, and $z$ are called the prefix, substring, and suffix of $S$, respectively. Let $\mathsf{Prefix}(S)$, $\mathsf{Substr}(S)$, and $\mathsf{Suffix}(S)$ denote the sets of prefixes, substrings, and suffixes of $S$, respectively. The elements of $\mathsf{Prefix}(S)\setminus\{S\}$, $\mathsf{Substr}(S)\setminus\{S\}$, and $\mathsf{Suffix}(S)\setminus\{S\}$ are called the proper prefixes, proper substrings, and proper suffixes of $S$, respectively. For a string $S$ of length $n$, $S[i]$ denotes the $i$-th symbol of $S$ and $S[i..j]=S[i]\cdots S[j]$ denotes the substring of $S$ that begins at position $i$ and ends at position $j$ for $1\leq i\leq j\leq n$. For convenience, let $S[i..j]=\varepsilon$ for $i>j$. The reversed string of a string $S$ is denoted by $S^{R}$, that is, $S^{R}=S[|S|]\cdots S[1]$.
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, let $\mathsf{Substr}(\mathcal{S})=\bigcup_{i=1}^{k}\mathsf{Substr}(S_{i})$. Let $\|\mathcal{S}\|=\sum_{i=1}^{k}|S_{i}|$ denote the total length of the strings in $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$.
2.2 All-Pairs Suffix-Prefix Overlap (APSP) Problems
In this section, we introduce the problems that we tackle in this paper.
Let $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ be a set of non-empty strings, where $S_{i}\in\Sigma^{+}$ for $1\leq i\leq k$. For each ordered pair $\langle S_{i},S_{j}\rangle$ of two strings in $\mathcal{S}$, the longest string in the set $\mathsf{Suffix}(S_{i})\cap\mathsf{Prefix}(S_{j})$ is called the longest suffix-prefix overlap of $S_{i}$ and $S_{j}$, and is denoted by $\mathsf{lspo}(i,j)$. We encode each $\mathsf{lspo}(i,j)$ by a tuple $(i,j,|\mathsf{lspo}(i,j)|)$ using $O(1)$ space.
2.2.1 APSP Problems on Static Sets
Problem 1 (static-APSP).
The static all-pairs suffix-prefix overlaps (static-APSP) problem is, given a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ non-empty strings, to compute
$\mathcal{L_{S}}=\{\mathsf{lspo}(i,j)\mid 1\leq i\leq k,1\leq j\leq k\}.$
Note that the output size of Problem 1 is $|\mathcal{L_{S}}|=\Theta(k^{2})$ as we compute $\mathsf{lspo}(i,j)$ for all ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$.
Problem 2 (length-bounded static-APSP).
The length-bounded static-APSP problem is, given a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ non-empty strings and an integer threshold $\ell\geq 0$, compute
$\mathcal{L_{S}^{\geq\ell}}=\{\mathsf{lspo}(i,j)\mid 1\leq i\leq k,1\leq j\leq k,|\mathsf{lspo}(i,j)|\geq\ell\}.$
2.2.2 APSP on Dynamic Sets with Insertions
We consider a dynamic variant of APSP allowing for insertions of strings, named dynamic APSP, that is defined as follows:
Problem 3 (dynamic APSP).
The dynamic APSP problem is, each time a new string $S_{i}$ is inserted to the current set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute the two following sets
$\mathcal{F}_{i}$ $=$ $\{\mathsf{lspo}(i,j)\mid 1\leq j\leq i\},$
$\mathcal{B}_{i}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i\}.$
We remark that $\mathcal{L_{S}}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}\cup\mathcal{B}_{i}\right)$ holds for the set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings. Thus, a solution to Problem 5 is also a solution to Problem 1.
Problem 4 (length-bounded dynamic APSP).
Let $\ell>0$ be a length threshold. The dynamic-APSP problem is, each time a new string $S_{i}$ is inserted to the current set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute the two following sets
$\mathcal{F}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(i,j)\mid 1\leq j\leq i,|\mathsf{lspo}(i,j)|\geq\ell\},$
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(j,i)|\geq\ell\}.$
Analogously, $\mathcal{L_{S}}^{\geq\ell}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}^{\geq\ell}\cup\mathcal{B}_{i}^{\geq\ell}\right)$ holds for the set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings. Thus, a solution to Problem 6 is also a solution to Problem 2 as well.
2.2.3 APSP on Fully Dynamic Sets with Insertions and Deletions
We also consider a fully dynamic variant of APSP allowing for insertions and deletions of strings, named fully dynamic APSP, that is defined as follows:
Problem 5 (fully dynamic APSP).
In the fully dynamic APSP problem, when a new string $S_{i}$ is added to the current set $\mathcal{S}$ of strings, the goal is to compute the two following sets
$\mathcal{F}_{i}$ $=$ $\{\mathsf{lspo}(i,j)\mid j\in\mathsf{ind}(\mathcal{S})\},$
$\mathcal{B}_{i}$ $=$ $\{\mathsf{lspo}(j,i)\mid j\in\mathsf{ind}(\mathcal{S})\},$
and add them to the list $\mathcal{L}_{\mathcal{S}}$, where $\mathsf{ind}(\mathcal{S})$ denotes the set of indices $h$ such that $S_{h}\in\mathcal{S}$. When a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings, then the goal is to remove $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ from the list $\mathcal{L}_{\mathcal{S}}$.
Problem 6 (length-bounded fully dynamic APSP).
Let $\ell>0$ be a length threshold. In the length-bounded fully dynamic APSP problem, when a new string $S_{i}$ is inserted to the current set $\mathcal{S}$ of strings, to compute the two following sets
$\mathcal{F}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(i,j)\mid 1\leq j\leq i,|\mathsf{lspo}(i,j)|\geq\ell\},$
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(j,i)|\geq\ell\},$
and add them to the list $\mathcal{L}_{\mathcal{S}}^{\geq\ell}$, where $\mathsf{ind}(\mathcal{S})$ denotes the set of indices $h$ such that $S_{h}\in\mathcal{S}$. When a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings, then the goal is to remove $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ from the list $\mathcal{L}_{\mathcal{S}}^{\geq\ell}$.
3 Tools
In this section we list the data structures that are used as building blocks of our static-APSP and dynamic-APSP algorithms.
3.1 Tries and Compact Tries
A trie $\mathsf{T}$ is a rooted tree $(\mathsf{V},\mathsf{E})$ such that
  • each edge in $\mathsf{E}$ is labeled by a single character from $\Sigma$ and
  • the character labels of the out-going edges of each node begin with distinct characters.
For a pair of nodes $u,v$ in a trie where $u$ is an ancestor of $v$, let $\langle u,v\rangle$ denote the path from $u$ to $v$. The path label of $\langle u,v\rangle$ is the concatenation of the edge labels from $u$ to $v$. The string label of a node $v$ in a trie, denoted $\mathsf{str}(v)$, is the path label from the root to $v$. The string depth of a node $v$ in a trie is the length of the string label of $v$, that is $|\mathsf{str}(v)|$.
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, let $\mathsf{Trie}(\mathcal{S})$ denote the trie that represents the strings in $\mathcal{S}$. We assume that each of the $k$ nodes in $\mathsf{Trie}(\mathcal{S})$ representing $S_{i}$ stores a unique integer $i$ called the id.
We build and maintain a compacted trie $\mathsf{ComTrie}(\mathcal{S})$ that is obtained from $\mathsf{Trie}(\mathcal{S})$ by keeping
  • all the leaves and the (possibly non-branching) internal nodes that store id’s $1,\ldots,k$ for the $k$ strings $S_{1},\ldots,S_{k}$, and
  • the root and all the branching nodes.
All other other nodes, each having a single child, are removed and their corresponding edges are contracted into paths. For each node $v$ in $\mathsf{Trie}(\mathcal{S})$, let $\mathsf{des}(v)$ denote the shallowest descendant of $v$ that exists in $\mathsf{ComTrie}(\mathcal{S})$. Note that the number of nodes in $\mathsf{ComTrie}(\mathcal{S})$ is $O(k)$.
Figure 1: Illustrations of $\mathsf{AC}(\mathcal{S})$ (left) and $\mathsf{ComTrie}(\mathcal{S})$ (right) for the set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings. The bold solid arcs represent trie edges and the dashed arcs represent failure links. The nodes representing the strings in $\mathcal{S}$ are depicted by double-lined circles with the string id’s.
3.2 Aho-Corasick Automata
For each non-root node $v$ of $\mathsf{Trie}(\mathcal{S})$, define its failure link by $\mathsf{flink}(v)=u$ iff $\mathsf{str}(u)$ is the longest proper suffix of $\mathsf{str}(v)$ that can be spelled out from the root of $\mathsf{Trie}(\mathcal{S})$. In other words, $\mathsf{str}(u)$ is the longest prefix of some string(s) represented by $\mathsf{Trie}(\mathcal{S})$ that is also a proper suffix of $\mathsf{str}(v)$. The failure links also form a (reversed) tree, and let us denote it by $\mathsf{FLTree}(\mathcal{S})$.
The Aho-Corasick automaton [2] for a set $\mathcal{S}$ of strings, denoted $\mathsf{AC}(\mathcal{S})$, is a finite automaton with two kinds of transitions [^1] We do not use the output function in our algorithms.: the goto function that is represented by the edges of the trie $\mathsf{Trie}(\mathcal{S})$, and the failure function that is represented by the edges of the reversed trie $\mathsf{FLTree}(\mathcal{S})$. It is clear that the total number of nodes and edges of $\mathsf{AC}(\mathcal{S})$ is $O(n)$, where $n=\|\mathcal{S}\|$. See Fig. 1 for concrete examples of the AC-automaton and its trie, compact trie, and failure links.
Theorem 1 ([7, 2]).
For a set $\mathcal{S}$ of strings of total length $n$, $\mathsf{AC}(\mathcal{S})$ can be built
  • in $O(n)$ time and space for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma)$ time and $O(n)$ space for a general ordered alphabet of size $\sigma$.
3.3 Directed acyclic word graphs (DAWGs)
For a set $\mathcal{S}$ of strings, let $\mathsf{End\_Pos}_{\mathcal{S}}(w)$ denote the set of pairs of string ids and ending positions of all occurrences of a substring $w\in\mathsf{Substr}(\mathcal{S})$, that is,
$\mathsf{End\_Pos}_{\mathcal{S}}(w)=\{(i,j)\mid S_{i}[j-|w|+1..j]=w,1\leq i\leq k,1\leq j\leq|S_{i}|\}.$
We consider an equivalence relation $\equiv_{\mathcal{S}}$ of strings over $\Sigma$ w.r.t. $\mathcal{S}$ such that, for any two strings $w$ and $u$, $w\equiv_{\mathcal{S}}u$ iff $\mathsf{End\_Pos}_{\mathcal{S}}(w)=\mathsf{End\_Pos}_{\mathcal{S}}(u)$. For any string $x\in\Sigma^{*}$, let $[x]_{\mathcal{S}}$ denote the equivalence class for $x$ w.r.t. $\equiv_{\mathcal{S}}$.
The DAWG of a set $\mathcal{S}$ of strings, denoted $\mathsf{DAWG}(\mathcal{S})$, is an edge-labeled DAG $(V,E)$ such that
$V$ $=$ $\{[{x}]_{\mathcal{S}}\mid x\in\mathsf{Substr}(\mathcal{S})\},$
$E$ $=$ $\{([{x}]_{\mathcal{S}},b,[{xb}]_{\mathcal{S}})\mid x,xb\in\mathsf{Substr}(\mathcal{S}),b\in\Sigma\}.$
We also define the set $L$ of suffix links of $\mathsf{DAWG}(\mathcal{S})$ by
$L=\{([{ax}]_{\mathcal{S}},a,[{x}]_{\mathcal{S}})\mid x,ax\in\mathsf{Substr}(\mathcal{S}),a\in\Sigma,[{ax}]_{\mathcal{S}}\neq[{x}]_{\mathcal{S}}\}.$
Namely, two substrings $x$ and $y$ in $\mathsf{Substr}(\mathcal{S})$ are represented by the same node of $\mathsf{DAWG}(\mathcal{S})$ iff the ending positions of $x$ and $y$ in the strings of $\mathcal{S}$ are equal.
See Fig. 2 and 3 for a concrete example of the DAWG and its suffix links.
Figure 2: $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is $\mathsf{Trie}(\mathcal{S})$.
Figure 3: Illustration of the suffix links of $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}$ of strings as in Fig 2.
For a node $v\in V$ of $\mathsf{DAWG}(\mathcal{S})$, let $\mathsf{long}(v)$ denote the longest string represented by $v$ (i.e., $|\mathsf{long}(v)|$ is the length of the longest path from the source to $v$). Also, let $\mathsf{slink}(v)=u$ denote the suffix link from node $u$ to node $v$.
An edge $(u,a,v)\in E$ of $\mathsf{DAWG}(\mathcal{S})$ is called an primary edge if $|\mathsf{long}(u)|+1=|\mathsf{long}(v)|$, and it is called a secondary edge otherwise ($|\mathsf{long}(u)|+1<|\mathsf{long}(v)|$). The primary edges form a spanning tree of $\mathsf{DAWG}(\mathcal{S})$ which consists of the longest paths from the source to all the nodes, and let us denote this spanning tree by $\mathsf{LPTree}(\mathcal{S})$. By the definition of the equivalence class $[{\cdot}]_{\mathcal{S}}$, for each $S_{i}\in\mathcal{S}$ it holds that $S_{i}=\mathsf{long}([{S_{i}}]_{\mathcal{S}})$. Thus we have the following:
Observation 1.
For each $S_{i}\in\mathcal{S}$, the path that spells out $S_{i}$ from the source of $\mathsf{DAWG}(\mathcal{S})$ is a path of $\mathsf{LPTree}(\mathcal{S})$.
$\mathsf{DAWG}(\mathcal{S})$ is a partial DFA of size $O(n)$ that accepts $\mathsf{Substr}(\mathcal{S})$ [4, 5]. The following lemma permits us to efficiently update the DAWG each time a new string is inserted to the current set of strings:
Theorem 2 ([5]).
Suppose that $\mathsf{DAWG}(\mathcal{S}_{i-1})$ has been built for a set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, one can update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time.
3.4 Suffix trees
The suffix tree for a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, denoted $\mathsf{STree}(\mathcal{S})$, is a compacted trie such that
  • each edge is labeled by a non-empty substring in $\mathsf{Substr}(\mathcal{S})$;
  • the labels of the out-edges of each node begin with distinct characters;
  • every internal node is branching (i.e. has two or more children) or represents a suffix in $\mathsf{Suffix}(\mathcal{S})$;
  • every suffix in $\mathsf{Suffix}(\mathcal{S})$ is represented by a node.
For ease of explanations, we identify each node of $\mathsf{STree}(\mathcal{S})$ with the string that the node represents.
The suffix link of a node $v$ in $\mathsf{STree}(\mathcal{S})$ points to the node that represents the string $\mathsf{str}(v)[2..|\mathsf{str}(v)|]$.
See Fig. 4 and 5 for a concrete example of the suffix tree and its suffix links.
Figure 4: $\mathsf{STree}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is a compacted version of $\mathsf{Trie}(\mathcal{S})$.
Figure 5: Illustration of the suffix links of $\mathsf{STree}(\mathcal{S})$ for the same set $\mathcal{S}$ of strings as in Fig 4.
Each edge label $x$ of $\mathsf{STree}(\mathcal{S})$ is represented by a tuple $(i,b,e)$ such that $x=S_{i}[b..e]$. This way $\mathsf{STree}(\mathcal{S})$ can be stored in $O(n)$ space, where $n=\|S\|$
Our version of suffix trees is the so-called Weiner tree [27] in which the repeated suffixes with unique right-extensions are represented by non-branching internal nodes (For example, string $\mathrm{bab}\in\mathsf{Suffix}(\mathcal{S})$ in Fig. 4 is represented by a non-branching internal node). This version of the suffix tree can easily be obtained by first building the suffix tree for strings $S_{1}\$_{1},\ldots,S_{k}\$_{k}$ with unique end-markers \$$\$_{1},\ldots,\$_{k}$ (\$$\$_{i}\neq\$_{j}$) and then by removing all edges labeled by the end-markers \$$\$_{1},\ldots,\$_{k}$.
By adapting Weiner’s construction [27] or Ukkonen’s construction [26] both working in online manners, one can update the suffix tree efficiently when a new string is inserted to the current set of strings. This setting is called as the semi-dynamic online construction of suffix trees in the literature [24].
Theorem 3 ([27, 26, 24]).
Suppose that $\mathsf{STree}(\mathcal{S}_{i-1})$ has been built for a set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, one can update $\mathsf{STree}(\mathcal{S}_{i-1})$ to $\mathsf{STree}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time.
4 Algorithm for Static-APSP
In this section, we present a new, very simple algorithm for the static-APSP problem for a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $\|\mathcal{S}\|=n$.
Our main data structure is the AC-automaton $\mathsf{AC}(\mathcal{S})$. For each string $S_{i}\in\mathcal{S}$ (in an arbitrary order), set node $v$ to be the node that represents $S_{i}$, i.e. initially $\mathsf{str}(v)=S_{i}$. All the nodes of $\mathsf{ComTrie}(\mathcal{S})$ are initially unmarked. Our algorithm is shown in Algorithm 1. See also Fig. 6 for illustration.
{itembox}
[c]Algorithm 1: Compute $\mathsf{lspo}(i,j)$ for all $1\leq j\leq k$
  1. (1)
    Set $v$ to be the node that represents $S_{i}$ (i.e. $\mathsf{str}(v)=S_{i}$).
  2. (2)
    • Set $u\leftarrow\mathsf{des}(v)$.
    • Traverse the induced subtree $\mathsf{T}_{u}$ of $\mathsf{ComTrie}(\mathcal{S})$ that is rooted at $u$, and consists only of unmarked nodes (and thus we do not traverse the subtrees under the marked nodes in $T_{u}$).
    • Report $\mathsf{lspo}(i,j)=|\mathsf{str}(v)|$ for all id’s $j$ found in the induced subtree $\mathsf{T}_{u}$.
    • Mark node $u$ on $\mathsf{ComTrie}(\mathcal{S})$.
  3. (3)
    If $v$ is the root, finish the procedure for the given string $S_{i}$. Unmark all marked nodes.
  4. (4)
    If $v$ is not the root, take the failure link from $v$ on $\mathsf{Trie}(\mathcal{S})$, and set $v\leftarrow\mathsf{flink}(v)$. Continue Step (2).
Figure 6: Illustration for Algorithm 1 that solves the static-APSP problem, in which we have climbed up the failure links from the node with id $i$ to node $v$. We access $u=\mathsf{des}(v)$ in $\mathsf{ComTrie}(\mathcal{S})$, and traverse only the white subtree below $u$.
Theorem 4.
For a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $n$, our algorithm solves the static-APSP problem with $O(n)$ working space and
  • in $O(n+k^{2})$ time for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma+k^{2})$ time for a general alphabet of size $\sigma$.
Proof.
The correctness of our algorithm follows from the definitions of the $\mathsf{lspo}(\cdot,\cdot)$ function and the failure links of $\mathsf{AC}(\mathcal{S})$, and from our failure link traversal starting from the node $v$ representing $S_{i}$, in decreasing order of the node depths towards the root $r$.
For each string $S_{i}$, Step (1) clearly takes $O(1)$ time. The total number $l$ of nodes $v,\mathsf{flink}(v),\mathsf{flink}^{2}(v),\ldots,\mathsf{flink}^{l-1}(v)=r$ visited in the chain of failure links starting from node $v$ with $\mathsf{str}(v)=S_{i}$ is at most $|S_{i}|+1$. The size of $\mathsf{ComTrie}(\mathcal{S})$ is $O(k)$. As we process the nodes $v,\mathsf{flink}(v),\mathsf{flink}^{2}(v),\ldots,\mathsf{flink}^{l-1}(v)=r$ in decreasing order of their depths, any marked node is visited at most once. This is because, if a marked node $w$ is visited twice, then there must be two distinct ancestors $p,q$ of $w$ in $\mathsf{ComTrie}(\mathcal{S})$, such that $p$ is an ancestor of $q$. This however is a contradiction, since $q$ was marked before $p$, and our algorithm does not visit the subtree under $q$ when processing the induced tree $\mathsf{T}_{p}$. Therefore, each node is marked at most once, and each marked node is visited at most once. Thus, the traversals on $\mathsf{ComTrie}(\mathcal{S})$ with our marking operations takes time linear in its size $O(k)$. Overall, for each $S_{i}\in\mathcal{S}$, we can compute $\mathsf{lspo}(i,j)$ for all $1\leq i\leq k$ in optimal $O(|S_{i}|+i)$ time, which leads to an overall $O(n+k^{2})$ time for all the $k$ strings in $\mathcal{S}$.
We build $\mathsf{AC}(\mathcal{S})$ by Theorem 1. We can compute $\mathsf{ComTrie}(\mathcal{S})$ and $\mathsf{des}(u)=v$ for all nodes $v$ in $\mathsf{Trie}(\mathcal{S})$ in time linear in the size of $\mathsf{Trie}(\mathcal{S})$, which is $O(n)$. ∎
For the length-bounded APSP problem, we have the following:
Corollary 1.
For a static set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings of total length $n$, our algorithm solves the $-\ell$ static-APSP problem with $O(n)$ working space and
  • in $O(n+|\mathcal{L_{\mathcal{S}}^{\geq\ell}}|)$ time for an integer alphabet of size $\sigma=n^{O(1)}$;
  • in $O(n\log\sigma+|\mathcal{L_{\mathcal{S}}^{\geq\ell}}|)$ time for a general alphabet of size $\sigma$.
Our data structure can be reused for different values of the length-threshold $\ell$.
Proof.
It suffices for us to change Step (3) and Step (4) in Algorithm 1 as follows:
  1. (3)
    If $|\mathsf{str}(v)|\leq\ell$, finish the procedure for the given string $S_{i}$. Unmark all marked nodes.
  2. (4)
    If $|\mathsf{str}(v)|>\ell$, take the failure link from $v$ on $\mathsf{Trie}(\mathcal{S})$, and set $v\leftarrow\mathsf{flink}(v)$. Continue Step (2).
We can reuse our data structure for different values of $\ell$, as we only need to change the underlined conditions in the above. ∎
5 Algorithm for Dynamic APSP
In this section, we present our algorithm for solving the dynamic APSP problem: Given a new string $S_{i}$ that is inserted to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, our goal is to compute the outputs $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ of Problem 5.
In what follows, we will show the following:
Theorem 5.
There exists a DAWG-based data structure of $O(n_{i})$ space, which solves the dynamic APSP problem of computing $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
Corollary 2.
Let $\ell>0$ be a length threshold. There exists a DAWG-based data structure of $O(n_{i})$ space, which solves the length-bounded dynamic APSP problem of computing $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
We will focus on proving Theorem 5. Corollary 2 can be obtained from Theorem 5, in a similar manner to the static case in Section 4.
5.1 Computing $\mathcal{F}_{i}$
We use the DAWG data structure for the dynamic case. Consider a set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. By Observation 1, $\mathsf{Trie}(\mathcal{S}_{i-1})$ is an induced tree of $\mathsf{LPTree}(\mathcal{S}_{i-1})$ that consists only of the paths spelling out $S_{1},\ldots,S_{i-1}$ from the source of $\mathsf{DAWG}(\mathcal{S}_{i-1})$. Also, if $\mathsf{flink}(u)=v$ for a node $u$ of $\mathsf{Trie}(\mathcal{S}_{i-1})$, then there is a chain of suffix links $u,\mathsf{slink}(u),\ldots,v$ of length $k\geq 1$ from the node $u$ to the node $v$ in $\mathsf{DAWG}(\mathcal{S}_{i-1})$ (i.e. $\mathsf{slink}^{k}(u)=v$). It is known that the suffix links of $\mathsf{DAWG}(\mathcal{S}_{i-1})$ form an edge-reversed tree that is equivalent to the suffix tree [27] of the reversed strings [5, 24]. We denote this suffix link tree by $\mathsf{SLTree}(\mathcal{S})$, and will use $\mathsf{SLTree}(\mathcal{S})$ in the dynamic case, instead of the tree $\mathsf{FLTree}(\mathcal{S})$ of failure links in the static case.
Now, we design our algorithm for the dynamic case with the DAWG structure, using the idea from Section 4 for the static case with the AC-automaton (see also Algorithm 1). Suppose that we have built and maintained the aforementioned data structures for the dynamic set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, we update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time with Theorem 2. We have now updated $\mathsf{SLTree}(\mathcal{S}_{i-1})$ to $\mathsf{SLTree}(\mathcal{S}_{i})$ as well.
To perform equivalent procedures as in Algorithm 1, we climb up the suffix link path from the node $v$ with $\mathsf{str}(v)=S_{i}$ toward the root on $\mathsf{SLTree}(\mathcal{S}_{i})$. Namely, we use $\mathsf{slink}$ instead of $\mathsf{flink}$. Each time we visit a node $v$ that is contained in $\mathsf{Trie}(\mathcal{S}_{i})$, we access $u=\mathsf{des}(v)$ on $\mathsf{ComTrie}(\mathcal{S}_{i})$ and perform the same procedures as in Algorithm 1. This gives us $\mathsf{lspo}(i,j)$ for all $1\leq j\leq i$ in a total of $O(|S_{i}|\log\sigma+i)$ time, since $\mathsf{ComTrie}(\mathcal{S}_{i})$ is of size $O(i)$. We can determine whether a node $v$ is in $\mathsf{Trie}(\mathcal{S}_{i})$ or not in $O(1)$ time after a simple $O(|S_{i}|\log\sigma)$-time preprocessing per new string $S_{i}$: After inserting $S_{i}$ to the DAWG, we simply trace the path that spells out $S_{i}$ from the source on the DAWG.
What remains is how to update $\mathsf{ComTrie}(\mathcal{S}_{i-1})$ to $\mathsf{ComTrie}(\mathcal{S}_{i})$, and update the $\mathsf{des}$ function. We do this by naïvely inserting the new string $S_{i}$ to $\mathsf{ComTrie}(\mathcal{S}_{i-1})$, traversing the tree with $S_{i}$ from the root in $O(|S_{i}|\log\sigma)$ time. The $\mathsf{des}$ function needs to be updated when an edge $(v,u)$ on $\mathsf{ComTrie}(\mathcal{S}_{i-1})$ is split into two edges $(v,w)$ and $(w,u)$, where $w$ is the parent of the new leaf representing $S_{i}$ on $\mathsf{ComTrie}(\mathcal{S}_{i})$. Let $v=v_{1},\ldots,v_{x}=w$ be the nodes in the uncompacted trie $\mathsf{Trie}(\mathcal{S}_{i})$ that are implicit on the edge $(v,w)$ in the compacted trie $\mathsf{ComTrie}(\mathcal{S}_{i})$. We update $\mathsf{des}(v_{1})\leftarrow w,\ldots,\mathsf{des}(v_{x})\leftarrow w$. We set $\mathsf{des}(y)\leftarrow s$ for all new nodes $y$ of $\mathsf{Trie}(\mathcal{S}_{i})$, where $s$ is the new leaf representing $S_{i}$. These updates can be done in $O(|S_{i}|)$ time.
To summarize this subsection, we have proven the following:
Lemma 1.
There is a DAWG-based data structure of $O(n_{i})$ space, which computes $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
Remark 1.
Diptarama et al. [11] showed how to simulate the failure link tree of the AC-automaton with the suffix link tree of the DAWG that is augmented with a nearest marked ancestor (NMA) data structure [28]. It should be noted that NMA data structures are not required in our algorithm.
5.2 Computing $\mathcal{B}_{i}$
We can easily reduce the problem of computing $\mathcal{B}_{i}$ to computing $\mathcal{F}_{i}$ for the set $\mathcal{S}_{i}^{R}=\{S_{1}^{R},\ldots,S_{i}^{R}\}$ of reversed strings, where the roles of suffixes and prefixes are swapped. This immediately gives us the following:
Corollary 3.
There is a DAWG-based data structure of $O(n_{i})$ space, which computes $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
6 Algorithm for Fully Dynamic APSP
In this section, we present a suffix tree based algorithm for the fully dynamic APSP problem allowing for insertions and deletions of strings.
We make use of the following observation in our algorithm.
Observation 2.
The induced tree of $\mathsf{STree}(\mathcal{S})$ that consists only of the paths spelling out the strings in $\mathcal{S}$ is a compacted version of $\mathsf{Trie}(\mathcal{S})$. The set of nodes of $\mathsf{ComTrie}(\mathcal{S})$ is a subset of the nodes of this induced tree.
See Fig. 1 and Fig. 4 for a concrete example of this observation.
6.1 Insertions
In the case of insertions, one can use essentially the same algorithm from Section 5: When a new string $S_{i}$ is inserted to the current set $\mathcal{S}$, we first update the suffix tree in $O(|S_{i}|\log\sigma)$ time by Theorem 3. Then, we climb up the suffix link path from the node representing $S_{i}$ toward the root. By Observation 2, the $\mathsf{des}(\cdot)$ function to access the corresponding nodes in $\mathsf{ComTrie}(\mathcal{S})$ can be maintained analogously. In more detail, by Observation 2, a single edge in $\mathsf{ComTrie}(\mathcal{S})$ may be split into multiple edges in the compacted version of $\mathsf{Trie}(\mathcal{S})$. Thus, when an edge $(u,v)$ in the compacted version of $\mathsf{Trie}(\mathcal{S})$ is split into two edges $(u,w)$ and $(w,v)$ during the update of the suffix tree, then we simply set $\mathsf{des}(w)\leftarrow\mathsf{des}(v)$. These all give us how to compute $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma)$ time using $\mathsf{STree}(\mathcal{S})$ of $O(n)$ space, where $n=\|\mathcal{S}\|$, as in Section 5.1.
One can compute $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma)$ time and $O(n)$ space by using $\mathsf{STree}(\mathcal{S}^{R})$ for the set $\mathcal{S}^{R}$ of reversed strings, as in Section 5.2.
6.2 Deletions
We can also extend our suffix-tree based algorithm in the case where both insertions and deletions are supported.
Suppose that a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings. Let $S_{i}[r..|S_{i}|]$ be the longest suffix of $S_{i}$ that corresponds to a branching node in the suffix tree, and let $S_{i}[t..|S_{i}|]$ be the longest suffix of $S_{i}$ such that $S_{i}[t..|S_{i}|]=S_{j}$ for some $j\in\mathsf{ind}(\mathcal{S})\setminus\{i\}$. Let $t=\infty$ when there is no such $j$ exists. Let $p=\min\{r,t\}$.
For each $1\leq q<p$ in increasing order, we remove the nodes $v_{q}$ (leaves or non-branching internal nodes) that respectively represent the suffixes $S_{i}[q..|S_{i}|]$. By tracing the suffix link chain from $v_{1}$ to $v_{p}$, this can be done in $O(1)$ time for each suffix $S_{i}[q..|S_{i}|]$ that is removed from the tree. If the parent of $v_{q}$ becomes non-branching after the removal of $v_{q}$, then the parent node is removed and its in-edge and out-edge are merged into a single edge. This is the same procedure as in the suffix tree in the sliding window model [15, 23, 16], and the total time complexity is $O(p\log\sigma)\subseteq O(|S_{i}|\log\sigma)$ for this step. Now the tree topology of the suffix tree has been updated.
Recall that each edge label $x$ is represented by a tuple $(i,b,e)$ such that $x=S_{i}[b..e]$. When $S_{i}$ is removed from $\mathcal{S}$, we need to replace a tuple $(i,b,e)$ which refers to $S_{i}$ with some tuple $(i^{\prime},b^{\prime},e^{\prime})$ such that $i^{\prime}\in\mathsf{ind}(\mathcal{S})\setminus\{i\}$ and $x=S_{i^{\prime}}[b^{\prime}..e^{\prime}]$. This can be done in $O(1)$ time by adapting the leaf-pointer algorithm [16] originally proposed for the sliding suffix trees. Leonard et al. [16] showed that edge-label maintenance in the suffix tree reduces to maintaining a pointer from each node to an arbitrary leaf in its subtree. They showed how to maintain such pointers in worst-case $O(1)$ time per leaf deletion. By applying this technique, we can update edge labels of $\mathsf{STree}(\mathcal{S})$ in a total of $O(p)\subseteq O(|S_{i}|)$ time.
Analogously, we update $\mathsf{STree}(\mathcal{S}^{R})$ in $O(|S_{i}|\log\sigma)$ time for the deleted string $S_{i}$.
In this fully dynamic setting, we maintain the output list $\mathcal{L}_{\mathcal{S}}$ with two arrays $F$ and $B$ such that $F[i]$ stores $\mathsf{lspo}(i,j)$ for all $j\in\mathsf{ind}(\mathcal{S})$ and $B[i]$ stores $\mathsf{lspo}(j,i)$ for all $j\in\mathsf{ind}(\mathcal{S})$. When a string $S_{i}$ is deleted, then we simply remove all elements in $F[i]$ and in $B[i]$. This clearly takes $O(k)$ time, where $k=|\mathcal{S}|$. We maintain these two dynamic arrays $F$ and $B$ by a standard doubling/halving technique: When all the entries of $F$ and $B$ become full, then we rebuild two arrays with $2|F|=2|B|$ entries each. When the number of occupied entries in $F$ and in $B$ reaches $|F|/4=|B|/4$, then we rebuild two arrays with $|F|/2=|B|/2$ entries each, and refresh the string ids. This takes $O(k)$ amortized time per insertion/deletion of a string, where $k$ is the number of strings in the set $\mathcal{S}$.
Finally, we obtain the following results:
Theorem 6.
There exists a suffix-tree-based data structure of $O(n)$ space, which solves the fully dynamic APSP problem in $O(|S_{i}|\log\sigma+k)$ amortized time for each inserted/deleted string $S_{i}$, where $n=\|\mathcal{S}\|$ and $k=|\mathcal{S}|$.
Corollary 4.
Let $\ell>0$ be a length threshold. There exists a suffix-tree-based data structure of $O(n)$ space, which solves the fully dynamic APSP problem in $O(|S_{i}|\log\sigma+k)$ amortized time for each inserted/deleted string $S_{i}$, where $n=\|\mathcal{S}\|$ and $k=|\mathcal{S}|$.
7 Conclusions and Future Work
In this paper, we presented an efficient algorithm for the APSP problem for a dynamic set of strings $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of total length $n$. Our algorithm occupies $O(n)$ space and takes $O(|S_{i}|\log\sigma+k)$ time for each string $S_{i}$ inserted or deleted. The running time is near-optimal except for the $\log\sigma$ factor, that is indeed required if one maintains $\mathcal{S}$ in a tree or DAG based data structure for a general ordered alphabet of size $\sigma$.
It is interesting to see whether our (fully) dynamic algorithm can be extended to efficient construction/update of the HOG [6, 12, 22] for a (fully) dynamic set of strings.
Acknowledgements
This work was supported by JSPS KAKENHI Grant Number JP23K24808 (SI).